# Elliptic curves and Fourier coefficients of meromorphic modular forms

### v1

@@ 6.2. The $U_{p}$ action
Fix $n\in\mathbb{Z}^{+}$ and a prime $p$. It suffices to prove that if $p^{l}{\hskip 1.0pt}|{\hskip 1.0pt}n$, then $p^{(s-1)l}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}(F)$. Let $t=v_{p}(A)$ and we will prove that $t$ is the largest integer such that $p^{2t}{\hskip 1.0pt}|{\hskip 1.0pt}m$ and $(-1)^{s-1}{\hskip 1.0pt}m/p^{2t}\equiv 0,1\pmod{4}$. Indeed, $p^{2t}{\hskip 1.0pt}|{\hskip 1.0pt}m$ since $p^{2t}{\hskip 1.0pt}|{\hskip 1.0pt}A^{2}$ and $A^{2}{\hskip 1.0pt}|{\hskip 1.0pt}m$, and
$(-1)^{s-1}\frac{m}{p^{2t}}\;=\;\bigg(\frac{A}{p^{t}}\bigg)^{2}(-1)^{s-1}{\hskip 1.0pt}\frac{m}{A^{2}}\;\equiv\;0,1\pmod{4}$.
Suppose that $t^{\prime}>t$ with $p^{2t^{\prime}}{\hskip 1.0pt}|{\hskip 1.0pt}m$ and $(-1)^{s-1}{\hskip 1.0pt}m/p^{2t^{\prime}}\equiv 0,1\pmod{4}$. Then,
$(-1)^{s-1}\frac{m}{p^{2t^{\prime}}}\;=\;\bigg(\frac{Ap^{t^{\prime}-t}}{p^{t^{\prime}}}\bigg)^{2}(-1)^{s-1}{\hskip 1.0pt}\frac{m}{(Ap^{t^{\prime}-t})^{2}}.$
By the maximality of $A$, we must have
$(-1)^{s-1}\frac{m}{p^{2t^{\prime}}}\;\equiv\;0\pmod{4}\quad\text{ and }\quad\bigg(\frac{Ap^{t^{\prime}-t}}{p^{t^{\prime}}}\bigg)^{2}\;\equiv\;0\pmod{4}$,
so in particular $2{\hskip 1.0pt}|{\hskip 1.0pt}A$ and $p\neq 2$. This then implies that
$(-1)^{s-1}\frac{m}{(Ap^{t^{\prime}-t})^{2}}\;=\;(-1)^{s-1}\frac{m}{A}\cdot\frac{1}{p^{2t^{\prime}-2t}}\;\equiv\;0,1\pmod{4}$,
contradicting the maximality of $A$.
Now, by Proposition 6.6, if $p^{l}{\hskip 1.0pt}|{\hskip 1.0pt}n$, then $p^{(s-1)l}{\hskip 1.0pt}|{\hskip 1.0pt}p^{(s-1)t}a_{n}(F_{m})$, where $F_{m}=\mathcal{S}_{d_{0}}f_{s+1/2,m}$. In particular, if $p^{l}{\hskip 1.0pt}|{\hskip 1.0pt}n$, then
$p^{(s-1)l}{\hskip 1.0pt}|{\hskip 1.0pt}A^{s-1}a_{n}(F_{m})\;=\;a_{n}(F).$
The result then follows. ∎
@@
Let $s=k/2$. As $s$ is even, one can apply Proposition 6.2 with $d_{0}=1$ and $d=D$ to obtain that
$\big(\mathcal{S}_{1}f_{s+1/2,-D}\big)(z)\;=\;(-1)^{\frac{s}{2}}{\hskip 1.0pt}|D|^{-\frac{s}{2}}{\hskip 1.0pt}\mathrm{Tr}_{D,1}\big((\partial_{\tau}^{s-1}G_{2s})(z,\tau)\big).$
Since $d_{0}=1$, we have that $\chi(Q)=1$ for all $Q\in\mathcal{Q}_{D}$. Hence, for an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$,
(15) $\mathrm{Tr}_{D,1}(f)\;=\;\sum_{Q\in\mathcal{Q}_{D}}\frac{1}{w_{Q}}f(\alpha_{Q}).$
@@
(16) $\mathcal{Q}_{D}\;=\;\bigsqcup_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\frac{A}{A^{\prime}}\cdot\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}},$
where $\frac{A}{A^{\prime}}\cdot\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}=\{\frac{A}{A^{\prime}}\cdot Q\mid Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}\}$. One can thus further write Equation 15 as
@@
$\big(\mathcal{S}_{1}f_{s+1/2,-D}\big)(z)$ $\;=\;(-1)^{\frac{s}{2}}{\hskip 1.0pt}|D|^{-\frac{s}{2}}{\hskip 1.0pt}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}{\hskip 1.0pt}(\partial_{\tau}^{s-1}G_{2s})(z,\alpha_{Q})$
@@
$\mathcal{G}_{2s,A^{2}D_{0}}$ $\;=\;\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mu\big(\tfrac{A}{A^{\prime}}\big)(-1)^{\frac{s}{2}}{\hskip 1.0pt}|A^{\prime 2}D_{0}|^{\frac{s}{2}}\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}$
$\;=\;(-1)^{\frac{s}{2}}|D_{0}|^{\frac{s}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mu\big(\tfrac{A}{A^{\prime}}\big)A^{\prime s}{\hskip 1.0pt}\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}.$
As $s$ is even, one can apply Proposition 6.7 to obtain that each $A^{\prime s-1}{\hskip 1.0pt}\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}$ is (s1)(s-1)-magnetic. It thus follows that $|D_{0}|^{-\frac{k}{4}}\mathcal{G}_{k,D}=|D_{0}|^{-\frac{s}{2}}\mathcal{G}_{2s,A^{2}D_{0}}$ has integer Fourier coefficients and is $\frac{k-2}{2}$-magnetic. Part (3) of Theorem 6.1 now follows from Proposition 6.5.
@@
Let $s=k/2$ and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}^{+}$ and some fundamental discriminant $D_{0}<0$ as before. As $s$ is odd, one can apply Proposition 6.2 with $d_{0}=D_{0}$ and $d=A^{2}$ to obtain that
$\big(\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}\big)(z)\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}{\hskip 1.0pt}\mathrm{Tr}_{A^{2},D_{0}}\big((\partial_{\tau}^{s-1}G_{2s})(z,\tau)\big).$
In this case, by Equation 16 and Lemma 6.8, we have that for an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$,
$\mathrm{Tr}_{A^{2},D_{0}}(f)\;=\;\sum_{Q\in\mathcal{Q}_{D}}\frac{1}{w_{Q}}\chi(Q)f(\alpha_{Q})\;=\;\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}\bigg(\frac{D_{0}}{A/A^{\prime}}\bigg){\hskip 1.0pt}f(\alpha_{Q}).$
@@
$\big(\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}\big)(z)$ $\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}\bigg(\frac{D_{0}}{A/A^{\prime}}\bigg){\hskip 1.0pt}(\partial_{\tau}^{s-1}G_{2s})(z,\alpha_{Q})$
$\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\bigg(\frac{D_{0}}{A/A^{\prime}}\bigg)\mathcal{G}_{2s,A^{\prime 2}D_{0}}(z).$
@@
$A^{s}{\hskip 1.0pt}\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}\;=\;(-1)^{\frac{s+1}{2}}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\bigg(\frac{D_{0}}{A/A^{\prime}}\bigg)\mathcal{G}_{2s,A^{\prime 2}D_{0}}(z).$
Here the left hand side is (s1)(s-1)-magnetic by Proposition 6.7. It follows from induction (on $A$) that $|D_{0}|^{\frac{k-2}{4}}\mathcal{G}_{k,D}=|D_{0}|^{\frac{s-1}{2}}\mathcal{G}_{2s,A^{2}D_{0}}$ has integer coefficients and is $\frac{k-2}{2}$-magnetic. Similarly, Part (3) of Theorem 6.1 follows from Proposition 6.5 and induction.
@@
$\chi_{A^{2},D_{0}}(Q)\;=\;\bigg(\frac{D_{0}}{A_{Q}\cdot p^{\prime}}\bigg)\;=\;\bigg(\frac{D_{0}}{A_{Q}}\bigg)\bigg(\frac{D_{0}}{p^{\prime}}\bigg),$

### v2

@@ 8.2. The Hecke actions
Fix a prime $p$. It suffices to prove that
$a_{np^{l}}(F|\lambda)\;\equiv\;0\pmod{p^{(s-1)l}}$
for all $n,l\in\mathbb{Z}^{+}$. Let $t=v_{p}(A)$ and let $m^{\prime}=m/p^{2t}$. Then, $m^{\prime}$ satisfies that $(-1)^{s-1}m^{\prime}\equiv 0,1\pmod{4}$. We will prove that the pair $(p,m^{\prime})$ satisfies the assumptions in Proposition 8.7.
Suppose that $p^{2}|m^{\prime}$. Then, $(Ap)^{2}|m$. If $p\geq 3$, then $(-1)^{s-1}m/(Ap)^{2}\equiv 0,1\pmod{4}$, contradicting the minimality of $A$. Hence, $p=2$. Write $A^{\prime}=A/p^{2t}$. By the maximality of $A$,
$(-1)^{s-1}\frac{m}{(2A)^{2}}\;=\;(-1)^{s-1}\frac{m^{\prime}}{4}\cdot\frac{1}{A^{\prime 2}}\;\equiv\;2,3\pmod{4}$.
Since $2\nmid A^{\prime}$, this implies that $(-1)^{s-1}m^{\prime}/4\equiv 2,3\pmod{4}$. Hence, the pair $(p,m^{\prime})$ satisfies the assumptions in Proposition 8.7.
Now, applying Proposition 8.7 with $m^{\prime}p^{2t}$ yields that
$a_{np^{l}}(F_{m^{\prime}p^{2t}}|\lambda)\;\equiv\;0\pmod{p^{(s-1)(l-t)}}$,
where $F_{m^{\prime}p^{2t}}=\mathcal{S}_{d_{0}}f_{s+1/2,m}$. The result then follows since $F=A^{s-1}F_{m^{\prime}p^{2t}}$ and $p^{t}|A$. ∎
@@
Adopt all the notations in Theorem 8.1. Assume that $s=k/2$ is even. As $s$ is even, one can apply Proposition 8.4 with $d_{0}=1$ and $d=D$ to obtain that
$\big(\mathcal{S}_{1}f_{s+1/2,-D}\big)(z)-h_{D}(z)\;=\;(-1)^{\frac{s}{2}}{\hskip 1.0pt}|D|^{-\frac{s}{2}}{\hskip 1.0pt}\mathrm{Tr}_{D,1}\big((\partial_{\tau}^{s-1}G_{2s})(z,\tau)\big),$
where $h_{D}:=h_{s,D,1}\in S_{2s}$ has integer coefficients by Lemma 8.3. Since $d_{0}=1$, we have that $\chi(Q)=1$ for all $Q\in\mathcal{Q}_{D}$. Hence, for an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$,
$\mathrm{Tr}_{D,1}(f)\;=\;\sum_{Q\in\mathcal{Q}_{D}}\frac{1}{w_{Q}}f(\alpha_{Q}).$ (30)
@@
$\mathcal{Q}_{D}\;=\;\bigsqcup_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\frac{A}{A^{\prime}}\cdot\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}},$ (31)
where $\frac{A}{A^{\prime}}\cdot\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}=\{\frac{A}{A^{\prime}}\cdot Q\mid Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}\}$. One can thus further write (30) as
@@
$\big(\mathcal{S}_{1}f_{s+1/2,-D}\big)(z)-h_{D}(z)$ $\;=\;(-1)^{\frac{s}{2}}{\hskip 1.0pt}|D|^{-\frac{s}{2}}{\hskip 1.0pt}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}{\hskip 1.0pt}(\partial_{\tau}^{s-1}G_{2s})(z,\alpha_{Q})$
@@
$\mathcal{G}_{2s,A^{2}D_{0}}$ $\;=\;\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mu\big(\tfrac{A}{A^{\prime}}\big)(-1)^{\frac{s}{2}}{\hskip 1.0pt}|A^{\prime 2}D_{0}|^{\frac{s}{2}}(\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}-h_{A^{\prime 2}D_{0}})$
$\;=\;(-1)^{\frac{s}{2}}|D_{0}|^{\frac{s}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mu\big(\tfrac{A}{A^{\prime}}\big)A^{\prime s}(\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}-h_{A^{\prime 2}D_{0}}).$
Now, each $\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}$ and $h_{A^{\prime 2}D_{0}}$ have integer coefficients, so it follows that $\widetilde{\mathcal{G}}_{k,D}=|D_{0}|^{-\frac{k}{4}}\mathcal{G}_{k,D}=|D_{0}|^{-\frac{s}{2}}\mathcal{G}_{2s,A^{2}D_{0}}$ has integer coefficients. Since each $h_{A^{\prime 2}D_{0}}\in S_{2s}$ and $\lambda$ annihilates $S_{2s}$, we obtain that
$\widetilde{\mathcal{G}}_{2s,A^{2}D_{0}}|\lambda\;=\;(-1)^{\frac{s}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\mu\big(\tfrac{A}{A^{\prime}}\big)A^{\prime s}{\hskip 1.0pt}\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}|\lambda.$
As $s$ is even, each $A^{\prime s-1}{\hskip 1.0pt}\mathcal{S}_{1}f_{s+1/2,-A^{\prime 2}D_{0}}|\lambda$ is (s1)(s-1)-magnetic by Proposition 8.8, so it follows that $\widetilde{\mathcal{G}}_{k,D}|\lambda$ is $\frac{k-2}{2}$-magnetic. Part (2) of Theorem 8.1 now follows from Proposition 8.6.
@@
Adopt all the notations in Theorem 8.1. Assume that $s=k/2$ is odd. Write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}^{+}$ and some fundamental discriminant $D_{0}<0$ as before. As $s$ is odd, one can apply Proposition 8.4 with $d_{0}=D_{0}$ and $d=A^{2}$ to obtain that
$\big(\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}\big)(z)-h_{D}(z)\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}{\hskip 1.0pt}\mathrm{Tr}_{A^{2},D_{0}}\big((\partial_{\tau}^{s-1}G_{2s})(z,\tau)\big),$
where $h_{D}:=h_{s,A^{2},D_{0}}\in S_{2s}$ has integer coefficients by Lemma 8.3. In this case, by (31) and Lemma 8.9, we have that for an $\mathrm{SL}_{2}(\mathbb{Z})$-invariant function $f(\tau)$ on $\mathcal{H}$,
$\mathrm{Tr}_{A^{2},D_{0}}(f)\;=\;\sum_{Q\in\mathcal{Q}_{D}}\frac{1}{w_{Q}}\chi(Q)f(\alpha_{Q})\;=\;\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\frac{2}{w_{A^{\prime 2}D_{0}}}\big(\tfrac{D_{0}}{A/A^{\prime}}\big){\hskip 1.0pt}f(\alpha_{Q}).$
@@
$\big(\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}\big)(z)-h_{D}(z)$ $\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\;\sum_{Q\in\mathcal{Q}_{A^{\prime 2}D_{0}}^{\mathrm{prim}}}\hskip-5.69046pt\frac{2}{w_{A^{\prime 2}D_{0}}}\big(\tfrac{D_{0}}{A/A^{\prime}}\big){\hskip 1.0pt}(\partial_{\tau}^{s-1}G_{2s})(z,\alpha_{Q})$
$\;=\;(-1)^{\frac{s+1}{2}}A^{-s}{\hskip 1.0pt}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\big(\tfrac{D_{0}}{A/A^{\prime}}\big)\mathcal{G}_{2s,A^{\prime 2}D_{0}}(z).$
@@
$A^{s}(\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}-h_{D})\;=\;(-1)^{\frac{s+1}{2}}|D_{0}|^{\frac{s-1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\big(\tfrac{D_{0}}{A/A^{\prime}}\big)\mathcal{G}_{2s,A^{\prime 2}D_{0}}.$
Now, both $\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}$ and $h_{D}$ have integer coefficients, so it follows from induction (on $A$) that $\widetilde{\mathcal{G}}_{k,D}=|D_{0}|^{\frac{k-2}{4}}\mathcal{G}_{k,D}=|D_{0}|^{\frac{s-1}{2}}\mathcal{G}_{2s,A^{2}D_{0}}$ has integer coefficients. Similarly as before, we obtain that
$A^{s}{\hskip 1.0pt}\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}|\lambda\;=\;(-1)^{\frac{s+1}{2}}\sum_{A^{\prime}{\hskip 1.0pt}|{\hskip 1.0pt}A}\big(\tfrac{D_{0}}{A/A^{\prime}}\big)\widetilde{\mathcal{G}}_{2s,A^{\prime 2}D_{0}}|\lambda.$
As $s$ is odd, $A^{s-1}\mathcal{S}_{D_{0}}f_{s+1/2,A^{2}}|\lambda$ is (s1)(s-1)-magnetic by Proposition 8.8, so it follows from induction that $\widetilde{\mathcal{G}}_{k,D}|\lambda$ is $\frac{k-2}{2}$-magnetic. Similarly, Part (2) of Theorem 8.1 follows from Proposition 8.6 and induction.
@@
$\chi_{A^{2},D_{0}}(Q)\;=\;\big(\tfrac{D_{0}}{A_{Q}\cdot p^{\prime}}\big)\;=\;\big(\tfrac{D_{0}}{A_{Q}}\big)\big(\tfrac{D_{0}}{p^{\prime}}\big),$